# ALMAGAL VIII. Cataloging Hierarchical Mass Structure from Cores to Clumps across the Galactic Disk

### v1

@@ Appendix A Clump Statistics For Each Morphological Category
In Table 3, we summarize the statistical quantities for the violin plot distributions shown in Section 4.2.
@@
Appendix B The $\Sigma_{\text{surr}}-M_{\text{core}}$ relation using clump temperature assumption
For the results we present in Section 5.2 we use the temperature model presented in Section 3.3 to calculate $\Sigma_{\text{surr}}$ for dendrogram structures surrounding cores from the ALMAGAL core catalog. To better understand how our temperature assumption affects our results, we decided to make this calculation using the clump-scale temperature, $T_{\text{clump}}$, as well. In Figure 9, we present the $\Sigma_{\text{surr}}-M_{\text{core}}$ relation using our original temperature model as well as the clump temperature assumption. In general, we find that $\Sigma_{\text{surr}}$ values increase systematically, which makes sense since the lower assumed temperatures from the clump-scale will result in higher mass estimates for the structures. In general, we find that the general trends still hold, and that the extents observed in $M_{\text{core}}$ vs. $\Sigma_{\text{surr}}$ still depend on the evolutionary stage of the hosting clump.
@@ Appendix B The $\Sigma_{\text{surr}}-M_{\text{core}}$ relation using clump temperature assumption
Figure 9: A comparison of the results in Figure 8 when using $\Sigma_{\text{surr}}$ values calculated using the temperature model presented in Section 3.3 the non-feathered images (left panels) and the feathered images (right panels). As in Figure 8, the contours indicate the shape of the distribution for the inner [90, 70, 40]% of the data. Marginal plots indicate the 1D kernel density estimates of the dendrogram structure surface density and core mass on the x- and y-axes, respectively.
Appendix C Mass Calculations Using Feathered Data
@@
$f=\frac{F_{\text{ALMA}}}{F_{\text{ATLASGAL}}}=\left(\frac{\nu_{\text{ALMA}}}{\nu_{\text{ATLASGAL}}}\right)^{\alpha_{\nu}}=\left(\frac{219}{345}\right)^{3.5}\approx 0.2$ (C1)
The ALMAGAL and ATLASGAL data are centered at frequencies of $\sim$ 219 GHz and $\sim$ 345 GHz, respectively. If we use Equation C1 (with $\beta=1.5,\alpha_{\nu}=3.5$), we acquire a scale factor of $f\sim$ 0.2, which is fairly small.
@@
$\frac{|M_{\text{feather}}-M_{\text{ALMA}}|}{M_{\text{feather}}}\times 100.$ (C2)
@@
Appendix D Structure differences in the near and far sample
As discussed in Section 3.2, the near and far sample distributions exhibit a systematic difference in mass sensitivity, with the far sample being less sensitive to lower mass structures. Since the dendrogram output is dependent on the mass sensitivity (see Section 3.2), we ultimately extract less low-mass structures from the far sample images, as seen in Figure 12. The difference in mass sensitivity can also explain why the near sample contains the largest structures, despite the fact that the linear size for the near sample field of view is generally smaller than that of the far sample.

### v2

@@ Appendix A Clump Statistics For Each Morphological Category
In Table 3, we summarize the statistical quantities for the violin plot distributions shown in Section IV.2.
@@
Appendix B Effect of dendrogram parameters on results
Table 4: A summary of the dendrogram parameters used for different versions of the catalog.
$f_{\text{min}}$ $\delta$ $f_{\text{peak}}$ $n_{\text{pix}}$
[…]
To determine how the results are affected by our choice of dendrogram parameters, we generated several new versions of the catalog. We list the parameters used for each version in Table 4. All versions of the catalog do not change the main results of our paper. In Figures 9 and 10, we present the KS test results and the $\Sigma_{\text{surr}}-M_{\text{core}}$ relationship for the version of the catalog that had the most significant impact on our results. This version has the following parameters: $f_{\text{min}}=3\sigma_{\text{rms}},\delta=3\sigma_{\text{rms}},f_{\text{peak}}=3\sigma_{\text{rms}},n_{\text{pix}}=0.5\times A_{\text{beam,pix}}$. The primary difference between Figures 6 and 9 is that the $p$-values for $d_{\text{clump}}$ and $R_{\text{GCR}}$ increase overall (excluding pairs that include the Empty subsample.) In Figure 10, we find no significant change in the relationship between $\Sigma_{\text{surr}}$ and $M_{\text{core}}$.
[…]
Refer to captionRefer to caption
[…]
Appendix C The $\Sigma_{\text{surr}}-M_{\text{core}}$ relation using clump temperature assumption
For the results we present in Section V.2 we use the temperature model presented in Section III.3 to calculate $\Sigma_{\text{surr}}$ for dendrogram structures surrounding cores from the ALMAGAL core catalog. To better understand how our temperature assumption affects our results, we decided to make this calculation using the clump-scale temperature, $T_{\text{clump}}$, as well. In Figure 11, we present the $\Sigma_{\text{surr}}-M_{\text{core}}$ relation using our original temperature model as well as the clump temperature assumption. In general, we find that $\Sigma_{\text{surr}}$ values increase systematically, which makes sense since the lower assumed temperatures from the clump-scale will result in higher mass estimates for the structures. In general, we find that the general trends still hold, and that the extents observed in $M_{\text{core}}$ vs. $\Sigma_{\text{surr}}$ still depend on the evolutionary stage of the hosting clump.
@@ Appendix C The $\Sigma_{\text{surr}}-M_{\text{core}}$ relation using clump temperature assumption
Figure 11: A comparison of the results in Figure 8 when using $\Sigma_{\text{surr}}$ values calculated using the temperature model presented in Section III.3 the non-feathered images (left panels) and the feathered images (right panels). As in Figure 8, the contours indicate the shape of the distribution for the inner [90, 70, 40]% of the data. Marginal plots indicate the 1D kernel density estimates of the dendrogram structure surface density and core mass on the x- and y-axes, respectively.
Appendix D Mass Calculations Using Feathered Data
@@
$f=\frac{F_{\text{ALMA}}}{F_{\text{ATLASGAL}}}=\left(\frac{\nu_{\text{ALMA}}}{\nu_{\text{ATLASGAL}}}\right)^{\alpha_{\nu}}=\left(\frac{219}{345}\right)^{3.5}\approx 0.2$ (D1)
The ALMAGAL and ATLASGAL data are centered at frequencies of $\sim$ 219 GHz and $\sim$ 345 GHz, respectively. If we use Equation D1 (with $\beta=1.5,\alpha_{\nu}=3.5$), we acquire a scale factor of $f\sim$ 0.2, which is fairly small.
@@
$\frac{|M_{\text{feather}}-M_{\text{ALMA}}|}{M_{\text{feather}}}\times 100.$ (D2)
@@
Appendix E Structure differences in the near and far sample
As discussed in Section III.2, the near and far sample distributions exhibit a systematic difference in mass sensitivity, with the far sample being less sensitive to lower mass structures. Since the dendrogram output is dependent on the mass sensitivity (see Section III.2), we ultimately extract less low-mass structures from the far sample images, as seen in Figure 14. The difference in mass sensitivity can also explain why the near sample contains the largest structures, despite the fact that the linear size for the near sample field of view is generally smaller than that of the far sample.